its a family of three the husband is 5years older than the wife and the son is half his mothers age ...altogether they sum up to 100. Question:: how ald is da husband ,,wife. And son !?!? Added together da answer must be hundred
Answer by Jynald-(6)

let w be the wife's age,
5 + w be the age of husband
and w/2 be the age of their son
solve for age of wife
100 = (5+w) + w + w/2
100 = 5 + 2w + w/2
200 = 10 + 4w + w
200-10 = 5w
190/5 = 5w/5
38 = w
solve for son's age:
w/2
38/2 = 19
solve for husband's age:
5 + w
5 + 38 = 43
checking:
husband's age plus son's age plus wife's age equals 100
43 + 19 + 38 = 100
100 = 100